# EVF-SAM: Early Vision-Language Fusion for Text-Prompted Segment Anything Model

### v2

@@ Abstract
Segment Anything Model (SAM) has attracted widespread attention for its superior interactive segmentation capabilities with visual prompts while lacking further exploration of text prompts. In this paper, we empirically investigate what text prompt encoders (e.g., CLIP or LLM) are good for adapting SAM for referring expression segmentation and introduce the Early Vision-language Fusion-based SAM (EVF-SAM). EVF-SAM is a simple yet effective referring segmentation method which exploits multimodal prompts (i.e., image and text) and comprises a pre-trained vision-language model to generate referring prompts and a SAM model for segmentation. Surprisingly, we observe that: (1) multimodal prompts and (2) vision-language models with early fusion (e.g., BEIT-3) are beneficial for prompting SAM for accurate referring segmentation. Our experiments show that the proposed EVF-SAM based on BEIT-3 can obtain state-of-the-art performance on RefCOCO/+/g for referring expression segmentation and demonstrate the superiority of prompting SAM with early vision-language fusion. In addition, the proposed EVF-SAM with 1.32B parameters achieves remarkably higher performance while reducing nearly 82% of parameters compared to previous SAM methods based on large multimodal models.
@@ 4.2 Implementation Details
Method Foundation Models w/ SAM? Training Data RefCOCO RefCOCO+ RefCOCOg AVG
@@ 4.3 Main Results
We mainly report the cIoU metric of RefCOCO-series benchmarks and compare our proposed EVF-SAM with recent state-of-the-art methods in Tab. II The upper part of Tab. II presents traditional methods based on text encoders. Despite their advantages in terms of fewer parameters and faster inference speeds, these methods either achieve less competitive results or require vast amounts of data due to their lack of integration with foundation models. The methods listed in the lower portion of Tab. II are based on Large Multimodal Models (LMMs), achieving state-of-the-art (SOTA) performance but require significant computational resources. Our EVF-SAM achieves the highest average cIoU score across all RES benchmarks, using only limited data and manageable computation costs. Specifically, our EVF-SAM achieves SOTA performance on RefCOCOg [46, 47], predicating a stronger capability for handling longer text prompts than previous LMM-based models, which is counter-intuitive while showing the great potential of vision-language models for understanding instructions. In addition, the early fusion between the input image and text prompts can generate more informative embeddings than independent encoders as discussed in Sec. 4. Meanwhile, EVF-SAM also achieves competitive performance on RefCOCO/+ [31, 45], demonstrating the effectiveness of the proposed formulation of EVF-SAM.
@@ 4.4 Ablation Study
5 Qualitative Results

### v3

@@ Abstract
Segment Anything Model (SAM) has attracted widespread attention for its superior interactive segmentation capabilities with visual prompts while lacking further exploration of text prompts. In this paper, we empirically investigate what text prompt encoders (e.g., CLIP or LLM) are good for adapting SAM for referring expression segmentation and introduce the Early Vision-language Fusion-based SAM (EVF-SAM). EVF-SAM is a simple yet effective referring segmentation method which exploits multimodal prompts (i.e., image and text) and comprises a pre-trained vision-language model to generate referring prompts and a SAM model for segmentation. Surprisingly, we observe that: (1) multimodal prompts and (2) vision-language models with early fusion (e.g., BEIT-3) are beneficial for prompting SAM for accurate referring segmentation. Our experiments show that the proposed EVF-SAM based on BEIT-3 can obtain state-of-the-art performance on RefCOCO/+/g for referring expression segmentation and demonstrate the superiority of prompting SAM with early vision-language fusion. In addition, the proposed EVF-SAM with 1.32B parameters achieves remarkably higher performance while reducing nearly 82% of parameters compared to previous SAM methods based on large multimodal models. The code, demo and models are available at https://github.com/hustvl/EVF-SAM.
@@ 4.2 Implementation Details
Method Foundation Models w/ SAM? Training Data RefCOCO () RefCOCO+ () RefCOCOg () AVG
@@ 4.3 Main Results
We mainly report the cIoU metric of RefCOCO-series benchmarks and compare our proposed EVF-SAM with recent state-of-the-art methods in Tab. II The upper part of Tab. II presents traditional methods based on text encoders. Despite their advantages in terms of fewer parameters and faster inference speeds, these methods either achieve less competitive results or require vast amounts of data due to their lack of integration with foundation models. The methods listed in the lower portion of Tab. II are based on Large Multimodal Models (LMMs), achieving state-of-the-art (SOTA) performance but require significant computational resources. Our EVF-SAM achieves the highest average cIoU score across all RES benchmarks, using only limited data and manageable computation costs. Specifically, our EVF-SAM achieves SOTA performance on RefCOCOg [46, 47], predicating a stronger capability for handling longer text prompts than previous LMM-based models, which is counter-intuitive while showing the great potential of vision-language models for understanding instructions. In addition, the early fusion between the input image and text prompts can generate more informative embeddings than independent encoders as discussed in Sec. 4. Furthermore, we incorporate the recently proposed SAM-2 [53] into EVF-SAM, and Tab. II shows that EVF-SAM-2 obtains significantly better performance on RefCOCO/+/g [31, 45], demonstrating the effectiveness of the proposed formulation of EVF-SAM.
@@ 4.4 Ablation Study
BEIT-3-Large SAM-2-L 898M 82.7 84.1 80.0 76.3 80.1 71.8 77.0 78.4 78.8
4.5 Discussions
To unveil how the multimodal encoder contributes to prompting SAM with texts, we visualize the attention maps between the [CLS] token (prompt embeddings) and the image tokens from the last layer of BEIT-3. As shown in Fig. 5, the attention maps focus on the target objects and are consistent with the input text prompts. The deep fusion of text and image embeddings leads to accurate region-text alignment. Consequently, the prompt embeddings contain abundant object-related information, including semantics and spatial localization, which is conducive to SAM achieving precise object segmentation.
Figure 5: Visualizations of Attention Maps in Multimodal Encoder. To unveil the effects of the Multimodal Encoder, we visualize the attention maps between the [CLS] token and image tokens in the last layer of BEIT-3-Large. Specifically, we sum up the attention maps from all heads.
4.6 Qualitative Results